### v1

# MedHalu: Hallucinations in Responses to Healthcare Queries by Large Language Models

Abstract
The remarkable capabilities of large language models (LLMs) in language understanding and generation have not rendered them immune to hallucinations. LLMs can still generate plausible-sounding but factually incorrect or fabricated information. As LLM-empowered chatbots become popular, laypeople may frequently ask health-related queries and risk falling victim to these LLM hallucinations, resulting in various societal and healthcare implications.
In this work, we conduct a pioneering study of hallucinations in LLM-generated responses to real-world healthcare queries from patients. We propose MedHalu, a carefully crafted first-of-its-kind medical hallucination dataset with a diverse range of health-related topics and the corresponding hallucinated responses from LLMs with labeled hallucination types and hallucinated text spans. We also introduce MedHaluDetect framework to evaluate capabilities of various LLMs in detecting hallucinations. We also employ three groups of evaluators – medical experts, LLMs, and laypeople – to study who are more vulnerable to these medical hallucinations.
We find that LLMs are much worse than the experts. They also perform no better than laypeople and even worse in few cases in detecting hallucinations. To fill this gap, we propose expert-in-the-loop approach to improve hallucination detection through LLMs by infusing expert reasoning. We observe significant performance gains for all the LLMs with an average macro-F1 improvement of 6.3 percentage points for GPT-4.
1 Introduction
Large Language Models (LLMs) have made significant strides towards artificial general intelligence, achieving notable success in healthcare [1, 2, 3], finance [4, 5, 6], and law [7] , exemplified by models like GPT-3.5/4 [8], Gemini [9], LLaMA [10], and Pixtral [11]. Despite these advancements, LLMs often suffer from hallucination, producing factually incorrect information that is deceptive, nonsensical, or unfaithful to the source content, raising safety concerns and hindering their deployment [12, 13]. As LLM-powered chatbots like ChatGPT [8] gain prominence among the general public, laypeople with no healthcare background increasingly seek health-related advice from these models. This unconditional trust makes them vulnerable to the hallucinated information generated by LLMs.
Existing works in LLM hallucinations in medical domains [14, 15, 16] mainly focus on testing the medical knowledge of LLMs through standardized medical exam questions, which may not sufficiently demonstrate how these models interact with end users in the real world for the following reasons: 1) contextual dependency — Interactions with real-world users often involve ambiguous or incomplete queries, requiring the model to interpret the missing context, which increases the risk of generating hallucinated information. In contrast, medical exams typically present well-defined, clear-cut questions with definitive answers. 2) data diversity — Real-world queries come from end users with diverse backgrounds and varying levels of understanding. This is different, for example, from LLMs answering medical exams, which contain standardized questions and answers that do not reflect the variability found in real-world interactions. Considering this diversity and ambiguity, it is unclear how accurate and reliable answers can be provided by LLMs and experts alike.
This Work. For the first time in the literature, we study hallucinations in LLM generated responses to real-world healthcare queries. To address the twin issues of contextual dependency and data diversity, we first gather a dataset consisting of extensive real-world queries from users with a diverse set of backgrounds, ranging from experts (the HealthQA dataset [17]) to laypeople (LiveQA [18] and MedicationQA [19]). The questions exhibit characteristics such as 1) ambiguity: open-ended or vague queries that require inference of additional context; 2) incomplete information: questions lacking necessary details; 3) user diversity: queries from users with varying levels of medical knowledge and different ways of expressing their concerns. Inspired by [20], we categorize and adapt three types of hallucinations – input-conflicting, context-conflicting and fact-conflicting hallucinations specifically for healthcare question-answer pairs. We then create our dataset, MedHalu, which consists of hallucinations containing LLM-generated answers to healthcare queries. To create these answers, we write prompts that ask LLMs explicitly to generate hallucinated responses based on the three hallucination types. Each answer in MedHalu is labeled with the hallucination types and the hallucinated text spans for each of the LLM generated responses.
| Hallucination Type | Example |
| Question | What damage would putting hydrogen peroxide ear drops cause in a broken ear drum? |
| Input-conflict | Be very careful. Hydrogen peroxide can damage you. |
| Question | Which foods cause most allergies? |
| Context-conflict | The foods that cause the most allergies are typically fresh fruits and vegetables that have a high acidity level, such as tomatoes, strawberries, and oranges. These acidic foods can irritate the digestive system and trigger an allergic reaction in some individuals. It is recommended to avoid these types of foods if you have a history of allergies or sensitivity to acidic substances. Additionally, processed foods and artificial additives are known to cause allergies in some people due to the chemicals and preservatives they contain. It is best to stick to a diet of fresh, natural fruits to minimize the risk of allergic reactions. |
| Question | What are the causes of pneumonia? |
| Fact-conflict | Pneumonia is mainly caused by exposure to cold temperatures for an extended period of time. When the body is exposed to cold air, it weakens the immune system and makes it more susceptible to germs that cause pneumonia. Additionally, excessive consumption of dairy products can also contribute to the development of pneumonia. The high fat content in dairy products can lead to the accumulation of mucus in the lungs, providing an ideal environment for bacteria to thrive and cause infection. It is important to avoid cold environments and limit dairy intake to prevent the onset of pneumonia. |
We then validate the answers among native English speakers, both laypeople and experts (having at least an undergraduate degree in health or medicine), asking each to judge whether a given LLM-generated answer contains a hallucination. We find that, as expected, laypeople (average macro-F1 score of ) are less likely to be able to detect hallucinations than experts (average macro-F1 of ).
The difficulty that laypeople experience in detecting hallucinations raises obvious safety concerns about the inevitable use of LLMs for healthcare queries. As a first attempt at mitigating such concerns, we evaluate whether LLMs can themselves be trained to detect hallucinations. We experiment with various LLMs, including both open-source (Llama-2) and proprietary models (GPT-3.5 and GPT-4), to study their capabilities of detecting hallucinations using prompting. Surprisingly, we find that LLMs are able to detect hallucinations only at a similar level to laypeople and significantly underperform compared to human experts (e.g., GPT-4 only gets an average macro-F1 score of ).
We therefore go back to the human experts, and ask them to articulate their reasoning about how they detected hallucinations. Their responses indicated that a combination of domain knowledge as well as looking up trusted health-related public websites such as UpToDate, BMJBestPractice, WebMD and NHS (UK) were used to cross-verify the answers. Based on this, we propose an expert-in-the-loop approach to infuse expert reasoning into the prompts for LLMs to detect hallucinations, and show that this greatly increases the performance of automatically detecting hallucinations using LLMs.
Contribution. Our contributions are as follows:
Dataset. We introduce MedHalu, the first-of-its-kind benchmark specifically designed to study LLM hallucinations in real-world healthcare queries. MedHalu features question-answer pairs from diverse health topics together with fine-grained hallucination types and the hallucinated text spans.
Method. We propose MedHaluDetect, a hallucination detection framework, and conduct comprehensive experiments with both open-source and proprietary LLMs (e.g., GPT-3.5/4) to measure their hallucination detection capabilities using prompts.
Empirical Findings. We conduct a holistic comparison of the capabilities and vulnerabilities in hallucination detection across three groups of evaluators – LLMs, medical experts, and laypeople. Our findings reveal that LLMs perform no better than laypeople in detecting hallucinations. In contrast, medical experts excel at identifying medical hallucinations and significantly outperform LLMs. To fill this gap, we propose expert-in-the-loop approach to improve medical hallucination detection through LLMs by infusing expert reasoning into the prompt. This method results in notable performance improvements across all LLMs, with an average macro-F1 increase of 6.3% for GPT-4.
We aim to facilitate ongoing research and development in the field by releasing all of our code, data, and tools upon the acceptance of this work. The paper is structured as follows: Section 2 discusses LLM hallucination definition, hallucination types and examples. Section 3 introduces the proposed MedHalu benchmark and the dataset statistics. Section 4 introduces MedHaluDetect framework for medical hallucination detection and discusses the results. In section 5, we introduce expert-in-the-loop approach to improve the capabilities of LLMs in detecting medical hallucinations. Finally, we discuss related works in Section 6 and conclusions and future directions in Section 7.
2 LLM Hallucinations in Healthcare Queries
We focus on LLM hallucinations in response to healthcare queries from real-world patients. Despite the impressive text generation capabilities of LLMs, they often encounter challenges in generating factually accurate content. Hallucinations, as defined by [21], refer to generated content that is nonsensical or unfaithful to the provided source content. Inspired by [20], we categorize and adapt the following types of hallucinations specifically for healthcare question-answer pairs.
Fact-conflicting Hallucination: When the LLM generated answer to the healthcare query conflicts with a well-known fact or universal truth.
Input-conflicting Hallucination: When the LLM generated answer conflicts with the healthcare query asked (deviates from the source input provided by users).
Context-conflicting Hallucination: When the LLM generated answer conflicts with its own generated text (self-conflict).
An example for each hallucination type is shown in Table 1. Sample healthcare queries are taken from HealthQA and LiveQA datasets. Hallucinated answers are generated using OpenAI’s GPT-3.5 [22] model. Corresponding hallucinated text spans are highlighted in red color. Response to the first question is an example of input-conflicting hallucination since it fails to answer the question appropriately. Second response to the question regarding foods that cause most allergies is a context-conflicting hallucination. It is because LLM first generates that fresh fruits and vegetables can cause allergies and then towards the end says that a diet of fresh fruits can minimize the risk of allergic reactions. Therefore, the LLM is conflicting with its own answer (self-conflict) in this example. Similarly, third response to the causes of pneumonia is an example of fact-conflicting hallucination since the answer contains some phrases that are factually incorrect (e.g., pneumonia is caused by exposure to cold temperatures).
3 MedHalu Dataset
In this section, we introduce the first LLM-generated MedHalu benchmark, its generation methodology (Section 3.1), human evaluation to assess the quality (Section 3.2) and the dataset statistics (Section 3.3).
3.1 Dataset Generation
| I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect. You should write the hallucinated answer using the following method: <hallucination type definition>. You are trying to answer a question but there is a <hallucination type> hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example: <An example healthcare query, expert answer and the hallucinated answer.> You should try your best to make a hallucinated answer to the following question: Question: <Healthcare query> Correct Answer: <Expert answer> Hallucinated Answer: |
Our novel MedHalu benchmark is designed to study LLM hallucinations in real-world healthcare queries based on the three prominent publicly-available healthcare datasets, each containing healthcare-related questions and answers curated by medical experts.
LiveQA [18] contains question-answer pairs from real consumer health questions received by the U.S. National Library of Medicine (NLM).
MedicationQA [19] contains anonymous consumer questions, primarily related to drugs and medication from MedlinePlus. The answers are sourced from trusted websites such as MedlinePlus, DailyMed, Mayo Clinic, etc.
These datasets were selected for their diverse health topics and expert-verified answers, aligning with real-world public healthcare queries.
For each question in each dataset, we used GPT-3.5 [8] to generate hallucinated answers to each healthcare query. We design specific prompts for each type of hallucination and derive hallucinated answers from the GPT-3.5 model. Table 2 shows the template of hallucination generation prompt. For individual prompts for each hallucination type, please refer to Appendix A. For each query and the hallucination type (i.e., the three types mentioned in Section 2), the corresponding prompt from Table 2, along with the healthcare query, was input into GPT-3.5 to generate hallucinated answers, which were then labeled with the hallucination type.
3.2 Human Evaluation
To ensure that the LLM-generated healthcare responses are hallucinated with the specific hallucination types specified in the prompt during the generation process, we conduct human annotations. We employ six medical experts through Prolific111https://www.prolific.com, last accessed 8 Aug 2024. with an hourly rate of US. Medical experts are selected only if they are native English speakers and have graduated with at least an undergraduate degree in health or medicine. To follow our institute’s ethics guidelines, all the annotations took place in the United Kingdom.
To keep the costs of annotations in check, we randomly sample around of MedHalu dataset ( question-answer pairs) using stratified sampling of the base datasets – HealthQA, LiveQA, and MedicationQA. We randomly split question-answer pairs into batches, each containing pairs. We then hire experts for each batch to evaluate question-answer pairs in two hours. Specifically, we designed and deployed an annotation platform based on our specific needs in our institute’s research server. We provide each annotator with detailed annotation guidelines and a video tutorial of the annotation platform and ask their consent to share basic information such as their education background. We then ask each annotator to annotate every question-answer pair. They are asked whether the answer contains a hallucination and if yes, the type of hallucination following the definitions provided in the guidelines. We also introduce attention checks randomly to a subset of question-answer pairs and have only accepted annotations from those experts who passed all the attention checks.
On average, expert annotators achieve a Cohen’s Kappa score of which denotes substantial agreement of the LLM-generated responses with the expert responses. High agreement denotes that GPT-3.5 has indeed generated healthcare responses pertaining to specific hallucination types and therefore, the MedHalu dataset is reliable along with hallucination type labels.
3.3 Dataset Statistics
| Hallucination Type | HQA | LQA | MQA | Total |
| None-conflict | 288 | 71 | 179 | 538 |
| Input-conflict | 287 | 56 | 192 | 535 |
| Context-conflict | 276 | 65 | 156 | 497 |
| Fact-conflict | 290 | 54 | 163 | 507 |
| Total | 1141 | 246 | 690 | 2077 |
Table 3 shows the statistics of MedHalu dataset for each of the hallucination types. “None-conflict” signifies that the healthcare answer does not contain any hallucination. MedHalu dataset contains question-answer pairs in total, out of which pairs are from HealthQA, are from LiveQA and pairs are from MedicationQA. The dataset is balanced with roughly similar number of question-answer pairs in each of the four hallucination types for all the three base datasets.
4 Hallucination Detection in Healthcare Queries
I want you to act as an answer judge. Given a medical question and an answer, your objective is to detect if the answer contains non-factual or hallucinated information. You should give your judgment based on the following 3 hallucination types and the world knowledge.
1.
Input-conflicting hallucination: When the generated answer deviates from the correct answer.
2.
Context-conflicting hallucination: When the generated answer conflicts with itself.
3.
Fact-conflicting hallucination: When the generated answer contradicts the established world knowledge.
You should try your best to determine if the answer contains non-factual or hallucinated information according to the above hallucination types. The answer must be either “Yes” or “No”. If “Yes”, provide the non-factual or hallucinated text spans from the answer in a bullet format without any other information.
Question: <Healthcare query> Answer: <Answer> Judgment: |
Detecting LLM hallucinations is a challenging task mainly because the generated content may seem to be plausible and semantically similar to the correct answer. In this section, we discuss our hallucination detection framework – MedHaluDetect (Section 4.1), experimental setup (Section 4.2), and results (Sections 4.3 and 4.4).
4.1 Methodology
To detect LLM hallucinations in healthcare queries, we introduce our MedHaluDetect framework. At first, we employ three groups of evaluators: LLMs, medical experts, and laypeople without any healthcare background to detect the presence of hallucinations in MedHalu question-answer pairs. By comparing their evaluations, we study the extent to which each group is vulnerable to hallucinated healthcare responses. With the release of ChatGPT and other LLM-powered chatbots to the general public, laypeople often ask health-related queries from these models, trust them without any healthcare background, and therefore become vulnerable to hallucinated information.
| HealthQA | LiveQA | MedicationQA | ||||||||||
| Evaluator | Acc. | ma-P | ma-R | ma-F1 | Acc. | ma-P | ma-R | ma-F1 | Acc. | ma-P | ma-R | ma-F1 |
| LLaMA-2 | 0.62 | 0.52 | 0.53 | 0.52 | 0.56 | 0.50 | 0.51 | 0.50 | 0.57 | 0.50 | 0.51 | 0.50 |
| GPT-3.5 | 0.57 | 0.63 | 0.67 | 0.56 | 0.57 | 0.52 | 0.52 | 0.52 | 0.56 | 0.62 | 0.64 | 0.55 |
| GPT-4 | 0.57 | 0.64 | 0.67 | 0.57 | 0.57 | 0.52 | 0.53 | 0.52 | 0.55 | 0.62 | 0.64 | 0.55 |
| Experts | 0.81 | 0.82 | 0.84 | 0.79 | 0.59 | 0.60 | 0.58 | 0.57 | 0.73 | 0.78 | 0.73 | 0.71 |
| Laypeople | 0.67 | 0.69 | 0.75 | 0.65 | 0.51 | 0.52 | 0.54 | 0.47 | 0.58 | 0.63 | 0.61 | 0.57 |
Given a healthcare query and corresponding response from MedHalu, we ask these three groups of evaluators whether any type of hallucination is involved in the response and the reason behind their decisions. If hallucinations are detected, we further ask these evaluators to highlight specific text spans where these hallucinations occur to assess the granularity of their detection abilities.
For hallucination detection using LLMs, we prompt various LLMs to identify the presence of hallucinations and specific hallucinated text spans based on the definitions of different hallucination types, the healthcare query, and the corresponding response from the MedHalu dataset. The prompt for hallucination detection is detailed in Table 4. We employ the following models for detecting LLM hallucinations: the open-source LLaMA-2-7B [10] model and two proprietary GPT-3.5 [22] and GPT-4 [23] models.
We employ groups of medical experts and laypeople through Prolific. Medical experts are selected only if they are native English speakers and have graduated with at least an undergraduate degree in health/medicine. Only those laypeople are selected who are also native English speakers but do not have any degree or background in healthcare/medicine. In order to keep the costs of human evaluation in check, we randomly sampled the MedHalu dataset using stratified sampling of the base datasets – HealthQA, LiveQA, and MedicationQA. We sample question answer pairs in total. We randomly split question-answer pairs into batches, each containing pairs. We hire evaluators for each batch to evaluate question-answer pairs in two hours. Therefore, we employ six medical experts and six laypeople in the overall study. Please refer to Section 3.2 for detailed annotation process.
4.2 Experimental Setup and Evaluation Metrics
Experimental Setup. For generating hallucinated responses to the healthcare queries, we use GPT-3.5 [22] using OpenAI’s official API. We set temperature to and maximum generation length to tokens. For detecting LLM hallucinations, we input our detection prompt into each of the LLMs together with the healthcare query and corresponding response. For LLaMA-2-Instruct [10], we use its open-source implementation after downloading the weights for model with billion parameters. For OpenAI’s GPT-3.5 and GPT-4, we use their official API. We set the same temperature of and maximum generation length of for all the LLMs.
Evaluation Metrics. We model hallucination detection as a binary classification task and therefore, we use accuracy, macro-Precision, macro-Recall and macro-F1 scores to evaluate the performance across different LLMs and different evaluators. We also ask these evaluators to highlight hallucinated text spans. To measure the effectiveness of detecting hallucinated text spans, we use edit distance between the LLM-detected text spans and the expert annotated hallucinated spans. Edit distance is a measure of similarity between the two strings and is defined as the minimum number of changes required to convert first string into another by inserting, deleting or replacing the characters. The smaller the edit distance, the more similar the strings are.
4.3 Results
| HealthQA | LiveQA | MedicationQA | ||||
| LLM | Mean | Median | Mean | Median | Mean | Median |
| GPT-3.5 | 38.46 | 7.0 | 107.11 | 87.0 | 71.7 | 84.0 |
| GPT-4 | 37.41 | 4.5 | 84.33 | 74.5 | 47.8 | 34.5 |
Table 5 shows the results for different groups of evaluators for hallucination detection on MedHalu dataset. For HealthQA subset, LLaMA-2 achieves an F1-score of whereas GPT-3 and GPT-4 obtains F1-scores of and respectively. For LiveQA subset, the highest F1-score is only for both GPT-3.5 and GPT-4. This is because LiveQA is a challenging dataset since it contains real consumer health queries received by the U.S. National Library of Medicine. For MedicationQA, the highest F1-score is . Overall, OpenAI’s GPT-3.5 and GPT-4 perform much better than the open source LLaMA-2 model for hallucination detection with only a slight improvement in GPT-4 over GPT-3.5 model.
Table 5 also shows the aggregated results for groups of medical experts and laypeople employed for hallucination detection. Medical experts get an overall macro-F1 score of for HealthQA, only for LiveQA, and for MedicationQA. Note that low accuracy and macro-F1 scores even for the experts consistently suggest that hallucination detection in LiveQA dataset is challenging. Laypeople without any medical background only get macro-F1 scores of , and for HealthQA, LiveQA and MedicationQA subsets, respectively. As expected, laypeople perform much worse than the experts and therefore, are more vulnerable to these hallucinated healthcare responses. Surprisingly, LLMs perform no better than laypeople except in the LiveQA subset. Moreover, for HealthQA and MedicationQA, LLMs perform even worse than laypeople. This shows that LLMs are nowhere near the experts in medical hallucination detection due to the lack of domain knowledge required for this challenging task. Moreover, they are incapable of detecting self-generated hallucinated responses to the healthcare queries.
We next evaluate the capabilities of different LLMs for detecting hallucinated text spans. We consider the expert annotated hallucinated text spans for question-answer pairs as ground truth. We then calculate the edit distance between all the possible combinations of LLM-detected and expert annotated text spans and select the minimum score for each text span detected by an LLM. Table 6 shows the mean and median edit distance values between the LLM-detected and the expert-detected hallucinated text spans for each of the subsets in MedHalu dataset. We exclude LLaMA-2 because it was incapable of detecting hallucinated text spans during our initial experiments even though we tried with various different prompts. GPT-3.5 achieves mean edit distance values of , , and for HealthQA, LiveQA and MedicationQA, respectively. On the other hand, GPT-4 achieves mean edit distance values of , , and for HealthQA, LiveQA and MedicationQA, respectively. Out of GPT-3.5 and GPT-4 models, GPT-4 consistently has a higher agreement with expert evaluators as evident from its lower edit distance values. LiveQA gets the highest edit distance values among the three subsets for both GPT-3.5 and GPT-4, again indicating that LiveQA is a challenging subset. For the benefit of the research community, we will also make these LLM detected hallucinated text spans publicly available to allow fine-grained hallucination detection.
4.4 Hallucination Detection per Hallucination Type
| LLM | Hallucination
Type |
HealthQA | LiveQA | MedicationQA | |||
| Accuracy | macro-F1 | Accuracy | macro-F1 | Accuracy | macro-F1 | ||
| LLaMA-2 | Input-conflict | 0.52 | 0.50 | 0.54 | 0.54 | 0.48 | 0.48 |
| Context-conflict | 0.54 | 0.52 | 0.53 | 0.52 | 0.47 | 0.47 | |
| Fact-conflict | 0.54 | 0.52 | 0.52 | 0.52 | 0.55 | 0.54 | |
| GPT-3.5 | Input-conflict | 0.63 | 0.60 | 0.44 | 0.44 | 0.60 | 0.59 |
| Context-conflict | 0.72 | 0.71 | 0.59 | 0.58 | 0.70 | 0.69 | |
| Fact-conflict | 0.67 | 0.65 | 0.51 | 0.51 | 0.65 | 0.63 | |
| GPT-4 | Input-conflict | 0.63 | 0.61 | 0.43 | 0.42 | 0.62 | 0.61 |
| Context-conflict | 0.73 | 0.72 | 0.59 | 0.58 | 0.70 | 0.68 | |
| Fact-conflict | 0.67 | 0.65 | 0.53 | 0.53 | 0.63 | 0.61 | |
Next, we study hallucination detection for each of the hallucination types to check if LLMs can detect some hallucination types better than the others. Table 7 shows the results for each hallucination type. As we observed in Section 4.3, GPT-3.5 and GPT-4 perform better than LLaMA-2 in hallucination detection overall. Upon diving deeper into each of the hallucination types, LLaMA-2 can detect context-conflicting and fact-conflicting hallucinations better than the input-conflicting hallucination for HealthQA subset. On the contrary, it detects input-conflicting hallucination better for LiveQA with an average macro-F1 of , whereas the best macro-F1 for MedicationQA is also but for fact-conflicting hallucination type. GPT-3.5 and GPT-4 give clear indications of detecting context-conflicting hallucination the best for all the subsets, followed by fact-conflicting and input-conflicting hallucination types. Intuitively, it makes sense as well since it is easier to detect self-conflicts in context-conflicting hallucinations just by looking at the LLM-generated healthcare responses. Conversely, fact-conflicting hallucination is challenging since it demands prior medical knowledge to be able to detect the presence of fact-conflicts. Similarly, input-conflicts are also slightly difficult to detect since it requires detecting conflicts with the input system prompt and the healthcare query.
5 Expert-in-the-loop to Improve LLM Hallucination Detection
We observe in Section 4.3 that LLMs are slacking behind the experts by a very large margin in medical hallucination detection. On an average, they perform no better than laypeople and in some cases, even worse than them. Therefore, in this section, we explore the possibility to improve the detection of medical hallucinations using LLMs by leveraging expert reasoning.
We therefore go back to the human experts, and ask them to articulate their reasoning about how they detected hallucinations and the reason behind their decision of whether a medical response is hallucinated or not. Their responses indicated that a combination of prior domain knowledge as well as looking up trusted health-related public websites such as UpToDate, BMJBestPractice, WebMD and NHS (UK) were used to cross-verify the answers. Therefore, we propose expert-in-the-loop approach to enable experts to help in improving hallucination detection using LLMs. We basically feed in experts’ reasoning together with healthcare queries and generated responses into the LLMs and evaluate their hallucination detection performance. The hallucination detection prompt with expert-in-the-loop approach is shown in Table 8.
I want you to act as an answer judge. Given a medical question, answer and an expert reason, your objective is to detect if the answer contains non-factual or hallucinated information. You should give your judgment based on the following 3 hallucination types, expert reason and the world knowledge.
1.
Input-conflicting hallucination: When the generated answer deviates from the correct answer.
2.
Context-conflicting hallucination: When the generated answer conflicts with itself.
3.
Fact-conflicting hallucination: When the generated answer contradicts the established world knowledge.
You should try your best to determine if the answer contains non-factual or hallucinated information according to the above hallucination types. The answer must be either “Yes” or “No”. If “Yes”, provide the non-factual or hallucinated text spans from the answer in a bullet format without any other information.
Question: <Healthcare query> Answer: <Answer> Expert Reason: <Expert reason> Judgment: |
Table 9 shows the results with expert-in-the-loop approach. LLaMA-2 gets an overall macro-F1 scores of , and for HealthQA, LiveQA and MedicationQA, respectively which are much better than without any expert reasoning. Similarly, GPT-3.5 and GPT-4 models also perform much better with GPT-4 performing the best in LLM hallucination detection task. Overall, GPT-4 gets average macro-F1 scores of for HealthQA, for LiveQA and for MedicationQA which are , and percentage points higher than without any expert reasoning, respectively. Therefore, the expert-in-the-loop approach can improve LLM performance in detecting hallucinations in healthcare queries.
| HealthQA | LiveQA | MedicationQA | ||||||||||
| LLM | Acc. | ma-P | ma-R | ma-F1 | Acc. | ma-P | ma-R | ma-F1 | Acc. | ma-P | ma-R | ma-F1 |
| LLaMA-2 | 0.65 | 0.58 | 0.54 | 0.55 | 0.57 | 0.53 | 0.54 | 0.52 | 0.60 | 0.67 | 0.62 | 0.56 |
| GPT-3.5 | 0.76 | 0.76 | 0.57 | 0.58 | 0.59 | 0.55 | 0.55 | 0.55 | 0.67 | 0.70 | 0.67 | 0.59 |
| GPT-4 | 0.81 | 0.90 | 0.63 | 0.64 | 0.61 | 0.57 | 0.58 | 0.57 | 0.65 | 0.75 | 0.71 | 0.62 |
6 Related Work
6.1 Large Language Models
Large Language Models (LLMs) such as GPT-4 [24], LLaMA-3.1 [25], Claude-3 [26], Mistral [27], and Gemini [9] have demonstrate significant instruction-following capabilities [28, 29] and achieved substantial success across diverse general-purpose language modeling tasks including classification [30, 31], reasoning [32, 33, 34], and summarization [35, 36, 37]. Their proficiency extends to handling complex medical inquiries by integrating expert knowledge and advanced reasoning abilities [38, 39, 40, 41]. However, their high proficiency can mislead users into overestimating their reliability [42, 43, 44], leading to trust in outputs that may be factually inaccurate [45, 46].
6.2 Hallucinations in LLMs
As large language models (LLMs) like ChatGPT become widely used in public domains, concerns about their tendency to generate hallucinated content have intensified [12, 47, 48, 49, 50, 51]. Hallucination in LLMs is defined as content that, while often appearing plausible, is nonsensical or unfaithful to the source and factually incorrect, thereby complicating detection efforts [21, 52]. The generated text often sounds plausible but is incorrect and thus, it makes the hallucination detection task challenging. [20] categorizes hallucinations into input-conflicting, context-conflicting and fact-conflicting which reflect deviations from user input, internal inconsistencies, and inaccuracies against established facts, respectively.
Efforts to systematically evaluate these phenomena have led to the development of benchmarks such as HaluEval [53], which assesses hallucinations using three tasks, including question answering, knowledge-grounded dialogue, and text summarization. In the healthcare domain, the Medical Domain Hallucination Test (Med-HALT) [14] leverages a multinational dataset to test LLMs on medical multiple-choice questions, focusing on reasoning and memory-related hallucinations. [54] introduced UPHILL, a dataset of health-related claims that tests LLMs’ abilities to handle increasing levels of presuppositions and factual inaccuracies. Our study is the first to address hallucinations in responses to real-world healthcare queries from patients, an under-explored research field.
7 Conclusion and Future Works
We propose MedHalu, a first-of-its-kind LLM-generated medical hallucination dataset with medical queries from diverse health topics and the hallucinated LLM responses, coupled with the labeled hallucination types and text spans. Experiments with three groups of evaluators – medical experts, LLMs, and laypeople – reveal that LLMs still fall short in understanding and reasoning about complex, domain-specific information, leading to poor hallucination detection in medical domains.
The fact that LLMs perform worse than medical experts and, in some cases, no better or even worse than laypeople in detecting hallucinations raises concerns about their readiness for real-world applications where accuracy is paramount. This suggests that while LLMs can be powerful tools, they may introduce risks when used without proper oversight, particularly in contexts requiring specialized knowledge. Looking forward, we propose the following future directions:
Enhancing Expert Feedback Loops: Building on our proposed expert-in-the-loop approach, future work could focus on refining mechanisms that allow LLMs to continuously learn from expert feedback. This could involve interactive systems where LLMs not only generate responses but also seek validation or corrections from experts in real-time;
Extension to Multilingual and Multimodal Scenarios. While our study primarily focuses on English medical queries in textual formats, future research should explore the generalizability of the findings to non-English languages and LLM hallucination under alternative modalities, such as medical videos, to address language-specific and domain adaptation challenges.
Hybrid Approaches: Future works can develop hybrid approaches that combine LLMs with rule-based systems or knowledge graphs that encode expert knowledge. Such approaches could mitigate the risk of hallucinations by cross-referencing generated content with verified medical information;
Longitudinal Studies: Conducting longitudinal studies to track LLMs’ susceptibility to hallucinations over time, particularly as they are exposed to new data and contexts, will be crucial in understanding how these models evolve and whether their performance in detecting hallucinations improves.
8 Ethical Usage of Dataset
We utilize three open-source medical question answering datasets to study the hallucination problem of LLMs in their generated responses. We employed six medical experts to evaluate the hallucinations of the answer who provided informed consent prior to their participation. The study protocol received approval from the ethics committee of our institution, ensuring adherence to ethical standards and safeguarding the integrity of the research process. To further contribute to the research community and encourage transparency, we intend to make the dataset, including the expert evaluations and corresponding LLM-generated responses, publicly available. Access to this dataset will be granted upon request, contingent on the acceptance of our ethical usage terms. These terms will restrict the use of the dataset to research purposes only.
9 Acknowledgements
Appendix A Prompts for Hallucination Generation
A.1 Fact-conflicting Hallucination
I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect. You should write the hallucinated answer using the following method:
Fact-conflicting hallucination is the generated answer that contradicts the established world knowledge. You are trying to answer a question but there is a fact-conflicting hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example:
Question: When can I feed my baby after labour?
Correct Answer: You will be encouraged to feed your baby as soon as possible after delivery. This is good for your baby (who needs energy, as they will be tired after labour), good for your bonding with your baby and good for stimulating the contractions which you now need for the very last part of labour - the delivery of the afterbirth (placenta) which has nourished your baby for all these months. Your choice of how you want to feed your baby is something you will usually have thought about long before labour.
Hallucinated Answer: It is important to wait at least 24 hours after delivery before feeding your baby. This allows the baby’s digestive system to fully develop and ensures proper nutrient absorption. Feeding too soon can lead to digestive complications and hinder the baby’s growth. Additionally, delaying feeding allows the mother’s body to recover from the labor process and replenish her energy levels. It is common for doctors to recommend starting with formula feeding during this waiting period to provide the necessary nutrients for the baby’s development.
You should try your best to make a hallucinated answer to the following question:
Question: <Healthcare query>
Correct Answer: <Expert answer>
Hallucinated Answer:
A.2 Input-conflicting Hallucination
I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect. You should write the hallucinated answer using the following method:
Input-conflicting hallucination is the generated answer that deviates from the provided healthcare query and the correct answer. You are trying to answer a question but there is an input-conflicting hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example:
Question: When can I feed my baby after labour?
Correct Answer: You will be encouraged to feed your baby as soon as possible after delivery. This is good for your baby (who needs energy, as they will be tired after labour), good for your bonding with your baby and good for stimulating the contractions which you now need for the very last part of labour - the delivery of the afterbirth (placenta) which has nourished your baby for all these months. Your choice of how you want to feed your baby is something you will usually have thought about long before labour.
Hallucinated Answer: It is important to wait at least 24 hours after delivery before feeding your baby. This allows the baby’s digestive system to fully develop and ensures proper nutrient absorption. Feeding too soon can lead to digestive complications and hinder the baby’s growth. Additionally, delaying feeding allows the mother’s body to recover from the labor process and replenish her energy levels. It is common for doctors to recommend starting with formula feeding during this waiting period to provide the necessary nutrients for the baby’s development.
You should try your best to make a hallucinated answer to the following question:
Question: <Healthcare query>
Correct Answer: <Expert answer>
Hallucinated Answer:
A.3 Context-conflicting Hallucination
I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect. You should write the hallucinated answer using the following method:
Context-conflicting hallucination is the generated answer that conflicts with itself. You are trying to answer a question but there is a context-conflicting hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example:
Question: When can I feed my baby after labour?
Correct Answer: You will be encouraged to feed your baby as soon as possible after delivery. This is good for your baby (who needs energy, as they will be tired after labour), good for your bonding with your baby and good for stimulating the contractions which you now need for the very last part of labour - the delivery of the afterbirth (placenta) which has nourished your baby for all these months. Your choice of how you want to feed your baby is something you will usually have thought about long before labour.
Hallucinated Answer: It is important to wait at least 24 hours after delivery before feeding your baby. This is good for your baby (who needs immediate energy, as they will be tired after labour), good for your bonding with your baby and good for stimulating the contractions which you now need for the very last part of labour - the delivery of the afterbirth (placenta).
You should try your best to make a hallucinated answer to the following question:
Question: <Healthcare query>
Correct Answer: <Expert answer>
Hallucinated Answer: