### v2

# MedHalu: Hallucinations in Responses to Healthcare Queries by Large Language Models

Abstract
Large language models (LLMs) are starting to complement traditional information seeking mechanisms such as web search. LLM-powered chatbots like ChatGPT are gaining prominence among the general public. AI chatbots are also increasingly producing content on social media platforms. However, LLMs are also prone to hallucinations, generating plausible yet factually incorrect or fabricated information. This becomes a critical problem when laypeople start seeking information about sensitive issues such as healthcare. Existing works in LLM hallucinations in the medical domain mainly focus on testing the medical knowledge of LLMs through standardized medical exam questions which are often well-defined and clear-cut with definitive answers. However, these approaches may not fully capture how these LLMs perform during real-world interactions with patients.
This work conducts a pioneering study on hallucinations in LLM-generated responses to real-world healthcare queries from patients. We introduce MedHalu, a novel medical hallucination benchmark featuring diverse health-related topics and hallucinated responses from LLMs, with detailed annotation of the hallucination types and text spans. We also propose MedHaluDetect, a comprehensive framework for evaluating LLMs’ abilities to detect hallucinations. Furthermore, we study the vulnerability to medical hallucinations among three groups — medical experts, LLMs, and laypeople. Notably, LLMs significantly underperform human experts and, in some cases, even laypeople in detecting medical hallucinations. To improve hallucination detection, we propose an expert-in-the-loop approach that integrates expert reasoning into LLM inputs, significantly improving hallucination detection for all LLMs, including a macro-F1 improvement for GPT-4.
1 Introduction
Large Language Models (LLMs) have made significant strides towards artificial general intelligence, achieving notable success in domains such as healthcare (Cascella et al. 2023; Xu et al. 2024; Chen et al. 2024c), finance (Wu et al. 2023), and law (Cui et al. 2023), exemplified by models like GPT-4 (Achiam et al. 2023), GPT-3.5 (Ouyang et al. 2022), and LLaMA (Touvron et al. 2023). Despite these advancements, LLMs often suffer from hallucination, producing factually incorrect information that is deceptive, nonsensical, or unfaithful to the source content, raising safety concerns and hindering their deployment (Rawte, Sheth, and Das 2023).
For at least 10 years now, it has been recognised that people seek out answers to highly sensitive and personal medical information on the web and social media (De Choudhury, Morris, and White 2014). Researchers are starting to recognise that LLMs can complement traditional information seeking mechanisms like web search (Fernández-Pichel, Pichel, and Losada 2024). Indeed, as LLM answers start to get incorporated into Google search, the distinction between search and LLMs is starting to get blurred.
As LLM-powered chatbots like ChatGPT (OpenAI 2023a) gain prominence among the general public, laypeople with no healthcare background increasingly seek health-related advice from these models (Ayers et al. 2023). Even educated adults find it difficult to tell whether some information is generated by AI chatbots (KFF 2024). This unconditional trust makes them vulnerable to the hallucinated information generated by LLMs.
In this rapidly evolving information landscape, it becomes important to guard against the known pitfalls of LLMs such as hallucinations. Systematically measuring the dangers of hallucination in the real world is a hard problem. Existing works in LLM hallucinations in the medical domain (Pal, Umapathi, and Sankarasubbu 2023; Chen et al. 2024a; Vishwanath et al. 2024) focus mainly on testing the medical knowledge of LLMs through standardized medical exam questions. These approaches may not fully capture how these models perform in real-world interactions because of two important reasons: 1) Contextual Dependency. Real-world user queries can be ambiguous or incomplete, requiring the models to infer missing context, which increases the risk of generating hallucinated information. 2) User Variability. Real-world queries originate from end users with diverse backgrounds and varying levels of medical understanding. This is significantly different from LLMs answering standardized medical exams, which contain standardized well-defined, clear-cut questions with definitive answers that do not reflect the variability in real-world interactions. Thus, the accuracy and reliability of LLM responses in real-world scenarios remain an open question.
| Hallucination Type | Example |
| Question | What damage would putting hydrogen peroxide ear drops cause in a broken ear drum? |
| Input-conflict | Be very careful. Hydrogen peroxide can damage you. |
| Question | Which foods cause most allergies? |
| Context-conflict | The foods that cause the most allergies are typically fresh fruits and vegetables that have a high acidity level, such as tomatoes, strawberries, and oranges. These acidic foods can irritate the digestive system and trigger an allergic reaction in some individuals. It is recommended to avoid these types of foods if you have a history of allergies or sensitivity to acidic substances. Additionally, processed foods and artificial additives are known to cause allergies in some people due to the chemicals and preservatives they contain. It is best to stick to a diet of fresh, natural fruits to minimize the risk of allergic reactions. |
| Question | What are the causes of pneumonia? |
| Fact-conflict | Pneumonia is mainly caused by exposure to cold temperatures for an extended period of time. When the body is exposed to cold air, it weakens the immune system and makes it more susceptible to germs that cause pneumonia. Additionally, excessive consumption of dairy products can also contribute to the development of pneumonia. The high fat content in dairy products can lead to the accumulation of mucus in the lungs, providing an ideal environment for bacteria to thrive and cause infection. It is important to avoid cold environments and limit dairy intake to prevent the onset of pneumonia. |
This Work. We present the first study of LLM hallucinations in responses to real-world healthcare queries. To address the dual challenges of contextual dependency and user variability, we first collect a dataset of extensive real-world queries from users with varying medical expertise, ranging from healthcare professionals (HealthQA (Zhu et al. 2019)) to laypeople (LiveQA (Abacha et al. 2017) and MedicationQA (Abacha et al. 2019)). The questions feature: 1) ambiguity, where queries are open-ended or vague with multiple meanings, requiring inference of additional context; 2) incomplete information, where essential details are missing; and 3) user diversity, reflecting varying levels of medical knowledge.
Inspired by Zhang et al. (2023), we categorize hallucinations into three types—input-conflicting, context-conflicting, and fact-conflicting—for healthcare question-answer pairs. We then create MedHalu, a medical hallucination detection benchmark consisting of hallucinated LLM-generated answers to healthcare queries. Each answer in MedHalu is labeled with the hallucination type and the corresponding hallucinated text spans.
Insights. Our findings reveal a significant gap in hallucination detection between laypeople (macro-F1: ) and experts (macro-F1: ), highlighting safety concerns about relying on LLMs for healthcare advice.
We then ask whether LLMs can be recursively used to detect hallucinations. For example, if a layperson obtains an answer from an LLM to a sensitive healthcare query, we ask LLM, if that generated answer is provided as input to (the same or a different) LLM, whether this provided piece of text contains hallucinations.
We find that unfortunately LLMs are no better than laypeople. Indeed, they perform at a similar level (macro-F1: for GPT-4) to laypeople and significantly underperform human experts at detecting text containing hallucinations.
We then dig deeper and ask human experts what strategies they use to identify hallucination-containing text. We find that human experts rely on a combination of domain knowledge and trusted health resources such as UpToDate111https://www.wolterskluwer.com/en/solutions/uptodate, last accessed 9 Jan 2025., BMJBestPractice222https://bestpractice.bmj.com/info/us/, last accessed 9 Jan 2025., WebMD333https://www.webmd.com/, last accessed 9 Jan 2025., and NHS444https://www.nhs.uk/, last accessed 9 Jan 2025. to cross-validate the answers. Building on this, we propose an expert-in-the-loop approach that integrates expert reasoning into LLM prompts, significantly improving LLMs’ ability to automatically detect hallucinations.
I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect.
You should write the hallucinated answer using the following method: <hallucination type definition>. You are trying to answer a question but there is a <hallucination type> hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example: <An example healthcare query, expert answer and the hallucinated answer.> You should try your best to make a hallucinated answer to the following question: Question: <Healthcare query> Correct Answer: <Expert answer> Hallucinated Answer: |
Contribution. Our key contributions are:
Novel Dataset. We introduce MedHalu, the first medical hallucination detection benchmark specifically designed to study LLM hallucinations in real-world healthcare queries, featuring question-answer pairs from diverse health topics as well as fine-grained hallucination types and text spans.
Comprehensive Framework. We propose MedHaluDetect, a hallucination detection framework, and conduct evaluation across both open-source models and proprietary LLMs (e.g., GPT-3.5/4) to measure their detection capabilities.
Empirical Findings. We conduct a holistic comparison of the capabilities and vulnerabilities in hallucination detection across three groups of evaluators—LLMs, medical experts, and laypeople. Our findings reveal that LLMs perform no better than laypeople in detecting hallucinations. In contrast, medical experts excel at identifying medical hallucinations and significantly outperform LLMs.
Mitigation Strategy. To address this gap, we propose an expert-in-the-loop approach that integrates expert reasoning into LLM prompts, enhancing hallucination detection and resulting in improvements across all models and an average macro-F1 increase of % for GPT-4.
2 The MedHalu Benchmark
MedHalu is designed to study LLM hallucinations in real-world healthcare queries that are often asked by patients and laypeople on web forums. This section details the types of hallucinations we test for (Section 2.1), dataset generation process (Section 2.2), and human evaluation (Section 2.3).
2.1 Hallucination Types
Hallucinations occur when LLMs generate content that is nonsensical or unfaithful to the input (Ji et al. 2023). Inspired by Zhang et al. (2023), we identify three types of hallucinations. Examples for each type are shown in Table 1.
Fact-conflicting: the response contradicts a well-known fact or universal truth.
Input-conflicting: the response conflicts with or deviates from the input query.
Context-conflicting: the response is self-contradictory or internally inconsistent.
2.2 Dataset Generation
MedHalu is a medical hallucination detection benchmark based on the three publicly available, expert-curated healthcare datasets described below. Each dataset consists of open-ended healthcare-related questions and answers curated by medical experts.
HealthQA (Zhu et al. 2019) contains healthcare question-answer pairs constructed from healthcare articles on the popular health-services website Patient555https://patient.info, last accessed 9 Jan 2025.. The questions are created by medical experts from a diverse set of healthcare topics, and answers are sourced from the articles. It has an average (± standard deviation) query length of 7.72 ± 2.41 words. An example query is “What are the symptoms of tonsillitis?”.
LiveQA (Abacha et al. 2017) contains question-answer pairs from real consumer health questions received by the U.S. National Library of Medicine (NLM) with an average query length of 41.76 ± 37.38 words. An example query is “I work at the Airport and I am in contact with hundreds of people in a day. I want to be sure that I am not going to expose people to Shingles. When is it safe for me to go back to work?”.
MedicationQA (Abacha et al. 2019) contains anonymous consumer questions, primarily related to drugs and medication from MedlinePlus666https://medlineplus.gov, last accessed 9 Jan 2025. with an average query length of 6.86 ± 2.83 words. The answers are sourced from trusted websites such as MedlinePlus, DailyMed, Mayo Clinic, etc. An example query is “How many mg does it take to overdose on Oxycodone?”.
We select these three different subsets to have healthcare queries with diverse topics for hallucination dataset since the three datasets differ in terms of how these queries were obtained, the nature and the length of these queries. These queries align with real-world public healthcare queries. For each such healthcare query, we explicitly ask an LLM to generate hallucinations. Specifically, we design tailored prompts for each hallucination type (See Section 2.1) and feed them to GPT-3.5 (OpenAI 2023a) to capture answers that are expected to contain hallucinations.
Details of the hallucination generation prompts can be found in Table 2 and Appendix A. By carefully constructing these prompts, the MedHalu dataset captures a wide range of plausible healthcare-related hallucinations. This approach ensures that the generated examples are both realistic and relevant. While it may not encompass every possible real-world variation, we believe this provides an important starting point for assessing and mitigating hallucinations in the healthcare domain. We fully expect that future work will build on and improve on these, perhaps by adding additional kinds of queries or additional types of hallucinations.
2.3 Human Evaluation
To validate hallucinations in the LLM-generated responses, we employ medical experts through Prolific777https://www.prolific.com, last accessed 9 Jan 2025. with an hourly rate of US$18. Selected annotators were required to be native English speakers based in the UK with at least an undergraduate degree in health or medicine.
| Background | No. of Experts | |
| Education | Undergrad in Medicine | 14 |
| Graduate in Medicine | 16 | |
| Gender | Male | 18 |
| Female | 12 | |
| Other | 0 | |
| Race | White | 12 |
| Asian | 11 | |
| Black | 5 | |
| Other | 2 | |
To keep a balance between the cost and the quality, we randomly sample about of MedHalu ( question-answer pairs) using stratified sampling from HealthQA, LiveQA, and MedicationQA. These were split into batches of pairs each, with three medical experts assigned to each batch to conduct the evaluation in two hours. We designed and deployed a custom annotation platform (details in Appendix B), provided detailed annotation guidelines and a video tutorial, and obtained consent from the annotators to collect basic information about their background. Table 3 provides details of these medical experts and their background. This study also received approval from the ethics committee of our institution. For each question-answer pair, the experts were asked whether the answer contains any hallucination and the type of hallucination following the definitions provided in the guidelines. Experts were also asked to mark the hallucinated text spans in the answers for fine-grained hallucination detection. We also implemented random attention checks, and only the experts passing all these checks have their annotations accepted. The expert annotators achieve an average Cohen’s Kappa score of , denoting almost perfect agreement between experts and LLM-generated responses. This confirms the reliability of MedHalu as the LLM has indeed generated healthcare responses pertaining to specific hallucination types. The statistics of MedHalu for each hallucination type is shown in Table 4.
| Hallucination | HQA | LQA | MQA | Total |
| None-conflict | 288 | 71 | 179 | 538 |
| Input-conflict | 287 | 56 | 192 | 535 |
| Context-conflict | 276 | 65 | 156 | 497 |
| Fact-conflict | 290 | 54 | 163 | 507 |
| Total | 1141 | 246 | 690 | 2077 |
3 Hallucination Detection in Healthcare Queries
I want you to act as an answer judge. Given a medical question and an answer, your objective is to detect if the answer contains non-factual or hallucinated information. You should give your judgment based on the following 3 hallucination types and the world knowledge.
1.
Input-conflicting hallucination: When the generated answer deviates from the correct answer.
2.
Context-conflicting hallucination: When the generated answer conflicts with itself.
3.
Fact-conflicting hallucination: When the generated answer contradicts the established world knowledge.
You should try your best to determine if the answer contains non-factual or hallucinated information according to the above hallucination types. The answer must be either “Yes” or “No”. If “Yes”, provide the non-factual or hallucinated text spans from the answer in a bullet format without any other information.
Question: <Healthcare query> Answer: <Answer> Judgment: |
Detecting LLM hallucinations is particularly challenging because the generated content may seem to be plausible and semantically similar to the correct answer. In this section, we discuss our hallucination detection framework—MedHaluDetect (Section 3.1), experimental setup (Section 3.2), and evaluation metrics (Section 3.3).
3.1 Methodology
| HealthQA | LiveQA | MedicationQA | ||||||||||
| Evaluator | Acc | ma-P | ma-R | ma-F1 | Acc | ma-P | ma-R | ma-F1 | Acc | ma-P | ma-R | ma-F1 |
| LLaMA-2 | 0.62 | 0.52 | 0.53 | 0.52 | 0.56 | 0.50 | 0.51 | 0.50 | 0.57 | 0.50 | 0.51 | 0.50 |
| GPT-3.5 | 0.57 | 0.63 | 0.67 | 0.56 | 0.57 | 0.52 | 0.52 | 0.52 | 0.56 | 0.62 | 0.64 | 0.55 |
| GPT-4 | 0.57 | 0.64 | 0.67 | 0.57 | 0.57 | 0.52 | 0.53 | 0.52 | 0.55 | 0.62 | 0.64 | 0.55 |
| Experts | 0.81 | 0.82 | 0.84 | 0.79 | 0.59 | 0.60 | 0.58 | 0.57 | 0.73 | 0.78 | 0.73 | 0.71 |
| Laypeople | 0.67 | 0.69 | 0.75 | 0.65 | 0.51 | 0.52 | 0.54 | 0.47 | 0.58 | 0.63 | 0.61 | 0.57 |
Our MedHaluDetect framework for detecting LLM hallucinations in healthcare queries leverages input from three groups of evaluators: LLMs, medical experts, and laypeople without healthcare expertise. Given a healthcare query and its corresponding response from MedHalu, each group assesses whether the response contains any types of hallucination and provides justifications for their decisions. When hallucinations are detected, we further ask these evaluators to highlight specific text spans where these hallucinations occur to assess the granularity of their detection abilities.
Hallucination detection using LLMs. We prompt various models to identify the presence of hallucinations and the corresponding text spans based on the definitions of different hallucination types, the healthcare query, and the corresponding response from MedHalu. We employ both open-source models such as LLaMA-2 (Touvron et al. 2023) and proprietary models like GPT-3.5 (Ouyang et al. 2022) and GPT-4 (Achiam et al. 2023). The prompt for hallucination detection is detailed in Table 5. By comparing evaluations from all groups, we study their varying susceptibility to hallucinated healthcare responses.
| HealthQA | LiveQA | MedicationQA | ||||
| LLM | Mean | Med | Mean | Med | Mean | Med |
| GPT-3.5 | 38.46 | 7.0 | 107.11 | 87.0 | 71.7 | 84.0 |
| GPT-4 | 37.41 | 4.5 | 84.33 | 74.5 | 47.8 | 34.5 |
| HealthQA | LiveQA | MedicationQA | |||||
| LLM | Hallucination | Acc | ma-F1 | Acc | ma-F1 | Acc | ma-F1 |
| LLaMA-2 | Input-conflict | 0.52 | 0.50 | 0.54 | 0.54 | 0.48 | 0.48 |
| Context-conflict | 0.54 | 0.52 | 0.53 | 0.52 | 0.47 | 0.47 | |
| Fact-conflict | 0.54 | 0.52 | 0.52 | 0.52 | 0.55 | 0.54 | |
| GPT-3.5 | Input-conflict | 0.63 | 0.60 | 0.44 | 0.44 | 0.60 | 0.59 |
| Context-conflict | 0.72 | 0.71 | 0.59 | 0.58 | 0.70 | 0.69 | |
| Fact-conflict | 0.67 | 0.65 | 0.51 | 0.51 | 0.65 | 0.63 | |
| GPT-4 | Input-conflict | 0.63 | 0.61 | 0.43 | 0.42 | 0.62 | 0.61 |
| Context-conflict | 0.73 | 0.72 | 0.59 | 0.58 | 0.70 | 0.68 | |
| Fact-conflict | 0.67 | 0.65 | 0.53 | 0.53 | 0.63 | 0.61 | |
Hallucination detection by Experts and Laypeople. We employ groups of medical experts and laypeople through Prolific. Medical experts are selected only if they are native English speakers based in the UK and have graduated with at least an undergraduate degree in health/medicine. Similarly, only those laypeople, who are native English speakers based in the UK but do not have any degree or background in healthcare/medicine, are selected. In order to keep the costs of human evaluation in check, we randomly sampled the MedHalu dataset using stratified sampling of the base datasets—HealthQA, LiveQA, and MedicationQA. We sample question answer pairs in total. We randomly split question-answer pairs into batches, each containing pairs. We hire evaluators for each batch to evaluate question-answer pairs in two hours. Therefore, we employ medical experts and laypeople in the overall study. The detailed annotation process is discussed in Section 2.3 and Appendix B.
3.2 Experimental Setup
For generating hallucinated responses to the healthcare queries, we use GPT-3.5 (Ouyang et al. 2022) using OpenAI’s official API. We set temperature to and maximum generation length to tokens. For detecting LLM hallucinations, we input our detection prompt into each of the LLMs together with the healthcare query and corresponding response. For LLaMA-2-Instruct (Touvron et al. 2023), we use its open-source implementation after downloading the weights for model with billion parameters. For OpenAI’s GPT-3.5 and GPT-4, we use their official API. We set the same temperature of and maximum generation length of for all the LLMs. Our primary goal in including models with different parameters is not to claim superiority of any one model but to provide context for interpreting results and testing the challenging MedHalu benchmark.
3.3 Evaluation Metrics
We model hallucination detection as a binary classification task and thus leverage accuracy, macro precision (ma-P), macro-Recall (ma-R), and macro-F1 scores (ma-F1) as the evaluation metrics. We also ask the evaluators to highlight text spans which contain hallucinations. To measure the effectiveness of detecting hallucinated text spans, we measure the edit distance between the LLM-detected text spans and the expert annotated hallucinated spans. Edit distance measures the minimum number of changes (insertion, deletion, or substitution of characters) required to convert one string into the other. Smaller edit distances indicate greater similarity between LLM detected text spans and expert annotations.
4 Results
4.1 Overall Results
Performances of LLMs. Table 6 shows the results of different evaluator groups in hallucination detection on MedHalu. For HealthQA subset, LLaMA-2 achieves an F1-score of whereas GPT-3.5/4 achieve higher scores of and , respectively. On the more challenging LiveQA dataset, which contains real consumer health queries received by the U.S. National Library of Medicine, GPT-3.5/4 both achieve F1-score of only . For MedicationQA, the highest F1-score is . Overall, the proprietary models GPT-3.5/4 significantly outperform the open source LLaMA-2 for hallucination detection, with GPT-4 showing only marginal improvement over GPT-3.5.
Performances of Human Experts and Laypeople. Medical experts achieve macro-F1 scores of for HealthQA, for LiveQA, and for MedicationQA. The consistently low accuracy and macro-F1 scores highlight the difficulty of hallucination detection in LiveQA even for trained professionals. As expected, laypeople perform much worse than the experts, achieving macro-F1 scores of only , and for the three datasets and therefore, are more vulnerable to these hallucinated healthcare responses. Surprisingly, LLMs perform no better than laypeople except in the LiveQA subset, indicating that LLMs struggle with hallucination detection on specialized domains due to the lack of domain knowledge and are even unable to detecting self-generated hallucinated responses to the healthcare queries.
Results for detecting hallucinated text spans. We next evaluate the capabilities of different LLMs for detecting hallucinated text spans. Correctly identifying these hallucinated text spans is important because: a) this can allow the LLMs to be improved to generate fewer hallucinations; b) it can help laypeople and experts ignore the hallucinated part of the text. We consider the expert annotated hallucinated text spans for question-answer pairs as ground truth. We then calculate the edit distance between all the possible combinations of LLM-detected and expert annotated text spans and select the minimum score for each ground-truth hallucinated text span. Table 7 shows the mean and median edit distance values between the LLM-detected and the expert-detected hallucinated text spans for each of the subsets in MedHalu dataset. We exclude LLaMA-2 because it was incapable of detecting hallucinated text spans during our initial experiments even though we tried with various different prompts. GPT-3.5 achieves mean edit distance values of , , and for HealthQA, LiveQA and MedicationQA, respectively. On the other hand, GPT-4 achieves mean edit distance values of , , and for HealthQA, LiveQA and MedicationQA, respectively. Out of GPT-3.5 and GPT-4 models, GPT-4 consistently has a higher agreement with expert evaluators as evident from its lower edit distance values. LiveQA gets the highest edit distance values among the three subsets for both GPT-3.5 and GPT-4, again indicating that LiveQA is a challenging dataset. For the benefit of the research community, we will also make these LLM detected hallucinated text spans publicly available to allow fine-grained hallucination detection.
4.2 Are some types of hallucination easier to detect than others?
Next, we study hallucination detection for each of the hallucination types (Section 2.1) to check if LLMs can detect some hallucination types better than the others. Table 8 shows the results for each hallucination type. As we observed in Section 4, GPT-3.5 and GPT-4 perform better than LLaMA-2 in hallucination detection overall. Upon diving deeper into each of the hallucination types, LLaMA-2 can detect context-conflicting and fact-conflicting hallucinations better than the input-conflicting hallucination for HealthQA subset. On the contrary, it detects input-conflicting hallucination better for LiveQA with an average macro-F1 of , whereas the best macro-F1 for MedicationQA is also but for fact-conflicting hallucination type. GPT-3.5 and GPT-4 give clear indications of detecting context-conflicting hallucination the best for all the subsets, followed by fact-conflicting and input-conflicting hallucination types. This also makes sense since it is easier to detect self-conflicts in context-conflicting hallucinations just by looking at the LLM-generated healthcare responses. Conversely, fact-conflicting hallucination is challenging since it demands prior medical knowledge to be able to detect the presence of fact-conflicts. Similarly, input-conflicts are also slightly difficult to detect since it requires detecting conflicts with the input system prompt and the healthcare query.
5 Expert-in-the-loop to Improve LLM Hallucination Detection
We observe in Section 4 that LLMs underperform compared to experts by a very large margin in medical hallucination detection. On average, they perform no better than laypeople and in some cases, even worse than them. Therefore, in this section, we explore the possibility to improve the detection of medical hallucinations using LLMs by leveraging insights on how experts reason about the presence of hallucinations.
We therefore go back to the human experts, and ask them to articulate their reasoning about how they detected hallucinations and the reason behind their decision of whether a medical response is hallucinated or not. Their responses indicate that a combination of prior domain knowledge as well as looking up trusted health-related public websites such as UpToDate, BMJBestPractice, WebMD and NHS (UK) were used to cross-verify the answers. Therefore, we propose expert-in-the-loop approach to enable experts to help in improving hallucination detection using LLMs. We feed in experts’ reasoning together with healthcare queries and generated responses into the LLMs and evaluate their hallucination detection performance. The hallucination detection prompt with expert-in-the-loop approach is shown in Table 9.
I want you to act as an answer judge. Given a medical question, answer, and an expert reason, you should detect if the answer contains non-factual or hallucinated information. You should give your judgment based on the following 3 hallucination types, expert reasons, and world knowledge.
1.
Input-conflicting hallucination: When the generated answer deviates from the correct answer.
2.
Context-conflicting hallucination: When the generated answer conflicts with itself.
3.
Fact-conflicting hallucination: When the generated answer contradicts the established world knowledge.
You should try your best to determine if the answer contains non-factual or hallucinated information according to the above hallucination types. The answer must be either Yes or No. If Yes, provide the non-factual or hallucinated text spans from the answer in a bullet format without any other information.
Question: <Healthcare query> Answer: <Answer> Expert Reason: <Expert reason> Judgment: |
| HealthQA | LiveQA | MedicationQA | ||||||||||
| LLM | Acc | ma-P | ma-R | ma-F1 | Acc | ma-P | ma-R | ma-F1 | Acc | ma-P | ma-R | ma-F1 |
| LLaMA-2 | 0.65 | 0.58 | 0.54 | 0.55 | 0.57 | 0.53 | 0.54 | 0.52 | 0.60 | 0.67 | 0.62 | 0.56 |
| GPT-3.5 | 0.76 | 0.76 | 0.57 | 0.58 | 0.59 | 0.55 | 0.55 | 0.55 | 0.67 | 0.70 | 0.67 | 0.59 |
| GPT-4 | 0.81 | 0.90 | 0.63 | 0.64 | 0.61 | 0.57 | 0.58 | 0.57 | 0.65 | 0.75 | 0.71 | 0.62 |
Table 10 shows the results with expert-in-the-loop approach. LLaMA-2 gets an overall macro-F1 scores of , and for HealthQA, LiveQA and MedicationQA, respectively which are much better than without any expert reasoning. Similarly, GPT-3.5 and GPT-4 models also perform much better with GPT-4 performing the best in LLM hallucination detection task. Overall, GPT-4 gets average macro-F1 scores of for HealthQA, for LiveQA and for MedicationQA which are , and percentage points higher than without any expert reasoning, respectively. Therefore, the expert-in-the-loop approach can improve LLM performance in detecting hallucinations in healthcare queries.
6 Related Work
6.1 Large Language Models
Large Language Models (LLMs) such as GPT-4 (OpenAI 2023b), LLaMA-3 (AI 2024), Claude-3 (Anthropic 2024), Mistral (Jiang et al. 2023), and Gemini (Team et al. 2023) have achieved substantial success across diverse general-purpose language modeling tasks including classification, reasoning, and summarization (Srivastava et al. 2023; Zhu et al. 2023; Liu et al. 2023; Jin et al. 2024a, b). Their proficiency extends to handling complex medical inquiries by integrating expert knowledge and advanced reasoning abilities (Nori et al. 2023; Singhal et al. 2023a, b; Liévin et al. 2023). However, their high proficiency can mislead users into overestimating their reliability, leading to trust in outputs that may be factually inaccurate (Chen et al. 2024b).
6.2 Hallucinations in LLMs
As LLMs become widely used in public domains, concerns about their tendency to generate hallucinated content have intensified (Rawte, Sheth, and Das 2023; Deng et al. 2024; Chen et al. 2024c). Hallucination in LLMs is defined as content that, while often appearing plausible, is nonsensical or unfaithful to the source and factually incorrect, thereby complicating detection efforts (Ji et al. 2023; Chen et al. 2023b, a; Liu et al. 2022; Xu, Jain, and Kankanhalli 2024; Zhao et al. 2024). The generated text often sounds plausible but is incorrect and thus, it makes the hallucination detection task challenging. Zhang et al. (2023) categorizes hallucinations into input-conflicting, context-conflicting and fact-conflicting which reflect deviations from user input, internal inconsistencies, and inaccuracies against established facts, respectively.
6.3 Benchmarks for LLM Hallucinations
Efforts to systematically evaluate LLM hallucinations have led to the development of generic benchmarks such as HaluEval (Li et al. 2023), which assesses hallucinations in general contexts using three tasks, including question answering, knowledge-grounded dialogue, and text summarization. In the healthcare domain, the Medical Domain Hallucination Test (Med-HALT) (Pal, Umapathi, and Sankarasubbu 2023) provides a multinational dataset of multiple choice questions derived from medical examinations across various countries, focusing on reasoning and memory-related hallucinations. Vishwanath et al. (2024) detects hallucinations in the generated summaries from Electronic Health Records. Kaur, Choudhury, and Pruthi (2023) introduce UPHILL, a dataset of health-related claims that tests LLMs’ abilities to handle increasing levels of presuppositions and factual inaccuracies. All of these benchmarks are very different from the context of a lay person asking an LLM their medical queries, which is a highly common use case, especially after the prevalence of ChatGPT and other chatbots. Therefore, we introduce MedHalu and present the first study to address LLM hallucinations in responses to real-world healthcare queries from layperson patients.
7 Conclusions and Future Works
We propose MedHalu, a pioneering hallucination detection benchmark featuring diverse healthcare queries and corresponding LLM responses, annotated with hallucination types and text spans. Evaluation on medical experts, LLMs, and laypeople highlights that the MedHalu benchmark is challenging and also reveals the current limitations of LLMs in detecting hallucinations, particularly in complex, domain-specific scenarios.
This is precisely the motivation of our study, to highlight this current limitation, and to create a strong benchmark that can drive forward future research to overcome such limitations (e.g., through knowledge graphs (as noted below), medical literature and expert reasoning)
Looking forward, we propose several key directions for future research:
Mitigating Hallucination through Adaptation. MedHalu offers a rich corpus for fine-grained LLM hallucination detection. Fine-tuning LLMs using parameter-efficient techniques, such as LoRA (Hu et al. 2022) and QLoRA (Dettmers et al. 2024), on MedHalu can improve their reliability in real-world healthcare queries. Moreover, combining LLMs with rule-based systems, knowledge graphs (Agrawal et al. 2024) such as uncertainty-aware knowledge graphs (Chen, Chen, and Yang 2024) and medical literature that encode expert knowledge can mitigate hallucination risks by cross-referencing responses with verified medical information.
Enhancing Expert Feedback Loops. Building on our proposed expert-in-the-loop approach, future work could focus on refining mechanisms that allow LLMs to continuously learn from expert feedback. This could involve interactive systems where LLMs not only generate responses but also seek validation or corrections from experts in real-time.
Extension to Multilingual and Multimodal Scenarios. While our study primarily focuses on English medical queries in textual formats, future research can explore how LLMs handle inaccurate information in non-English languages and LLM hallucination under alternative modalities (e.g. medical videos and broadcast).
Limitations
The proposed MedHalu dataset contains real-world healthcare queries in English only. Therefore, it is unknown how LLMs would hallucinate in case of healthcare queries in non-English languages. In the future, we would like to focus on non-English queries as well to study LLM hallucinations. One possible approach can be to directly translate English healthcare queries into non-English languages to curate a multilingual dataset. Moreover, we are also in discussion with doctors in India for future work to potentially explore creating a benchmark for Hindi, Bengali and Marathi, which are among the 3 most spoken Indian languages.
As existing LLMs continue to train on more and more datasets and new LLMs keep releasing, hallucination detection may become increasingly challenging. It is important to keep up with that pace and continuously evaluate their ability to generate hallucinated text in order to ensure their safety and reliability. As healthcare requires its own domain knowledge, future work can build an agentic AI that can incorporate knowledge from either knowledge graphs or from specific sources of high quality medical literature (e.g., WebMD888https://www.webmd.com/ or Mayo Clinic articles999https://www.mayoclinic.org/diseases-conditions from the US, NHS help articles in the UK101010https://www.nhs.uk/health-a-to-z/ and others cited by our human experts as the strategies they use themselves in Section 5) to create better hallucination detectors for the limited context of real-world medical queries. This could also become part of upcoming AI regulations – to make high quality answers and/or the addition of hallucination detection filters as a key step for LLM providers.
Ethical Considerations
Ethical Usage of Dataset
We utilize three open-source medical question answering datasets to study the hallucination problem of LLMs in their generated responses. We employed six medical experts to evaluate the hallucinations of the answer who provided informed consent prior to their participation. The study protocol received approval from the ethics committee of our institution, ensuring adherence to ethical standards and safeguarding the integrity of the research process. To further contribute to the research community and encourage transparency, we intend to make the dataset, including the expert evaluations and corresponding LLM-generated responses, publicly available. Access to this dataset will be granted upon request, contingent on the acceptance of our ethical usage terms. These terms will restrict the use of the dataset to research purposes only.
Longitudinal Studies. The fact that LLMs perform worse than medical experts and, in some cases, no better or even worse than laypeople in detecting hallucinations raises concerns about their readiness for real-world applications where accuracy is paramount. This suggests that while LLMs can be powerful tools, they may introduce risks when used without proper oversight, particularly in contexts requiring specialized knowledge. Conducting longitudinal studies to track LLMs’ susceptibility to hallucinations over time, particularly as they are exposed to new data and contexts, will be crucial in understanding how these models evolve and whether their performance in detecting hallucinations improves.
Appendix A Hallucination Generation Prompts
A.1 Fact-conflicting Hallucination
I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect. You should write the hallucinated answer using the following method:
Fact-conflicting hallucination is the generated answer that contradicts the established world knowledge. You are trying to answer a question but there is a fact-conflicting hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example:
Question: When can I feed my baby after labour?
Correct Answer: You will be encouraged to feed your baby as soon as possible after delivery. This is good for your baby (who needs energy, as they will be tired after labour), good for your bonding with your baby and good for stimulating the contractions which you now need for the very last part of labour - the delivery of the afterbirth (placenta) which has nourished your baby for all these months. Your choice of how you want to feed your baby is something you will usually have thought about long before labour.
Hallucinated Answer: It is important to wait at least 24 hours after delivery before feeding your baby. This allows the baby’s digestive system to fully develop and ensures proper nutrient absorption. Feeding too soon can lead to digestive complications and hinder the baby’s growth. Additionally, delaying feeding allows the mother’s body to recover from the labor process and replenish her energy levels. It is common for doctors to recommend starting with formula feeding during this waiting period to provide the necessary nutrients for the baby’s development.
You should try your best to make a hallucinated answer to the following question:
Question: <Healthcare query>
Correct Answer: <Expert answer>
Hallucinated Answer:
A.2 Input-conflicting Hallucination
I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect. You should write the hallucinated answer using the following method:
Input-conflicting hallucination is the generated answer that deviates from the provided healthcare query and the correct answer. You are trying to answer a question but there is an input-conflicting hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example:
Question: When can I feed my baby after labour?
Correct Answer: You will be encouraged to feed your baby as soon as possible after delivery. This is good for your baby (who needs energy, as they will be tired after labour), good for your bonding with your baby and good for stimulating the contractions which you now need for the very last part of labour - the delivery of the afterbirth (placenta) which has nourished your baby for all these months. Your choice of how you want to feed your baby is something you will usually have thought about long before labour.
Hallucinated Answer: It is important to wait at least 24 hours after delivery before feeding your baby. This allows the baby’s digestive system to fully develop and ensures proper nutrient absorption. Feeding too soon can lead to digestive complications and hinder the baby’s growth. Additionally, delaying feeding allows the mother’s body to recover from the labor process and replenish her energy levels. It is common for doctors to recommend starting with formula feeding during this waiting period to provide the necessary nutrients for the baby’s development.
You should try your best to make a hallucinated answer to the following question:
Question: <Healthcare query>
Correct Answer: <Expert answer>
Hallucinated Answer:
A.3 Context-conflicting Hallucination
I want you to act as a hallucination answer generator. Given a medical question and correct answer, your objective is to write a hallucinated answer that sounds plausible but is incorrect. You should write the hallucinated answer using the following method:
Context-conflicting hallucination is the generated answer that conflicts with itself. You are trying to answer a question but there is a context-conflicting hallucination in the generated answer. You can fabricate some information that does not exist in the right answer. Below is an example:
Question: When can I feed my baby after labour?
Correct Answer: You will be encouraged to feed your baby as soon as possible after delivery. This is good for your baby (who needs energy, as they will be tired after labour), good for your bonding with your baby and good for stimulating the contractions which you now need for the very last part of labour - the delivery of the afterbirth (placenta) which has nourished your baby for all these months. Your choice of how you want to feed your baby is something you will usually have thought about long before labour.
Hallucinated Answer: It is important to wait at least 24 hours after delivery before feeding your baby. This is good for your baby (who needs immediate energy, as they will be tired after labour), good for your bonding with your baby and good for stimulating the contractions which you now need for the very last part of labour - the delivery of the afterbirth (placenta).
You should try your best to make a hallucinated answer to the following question:
Question: <Healthcare query>
Correct Answer: <Expert answer>
Hallucinated Answer:
Appendix B Annotation Platform For Human Evaluation
For LLM hallucination detection, we hire two sets of human evaluators–medical experts and laypeople through Prolific. We develop a customized annotation platform for annotating LLM hallucinated responses to the healthcare queries. The screenshot of the annotation guidelines page is shown in Figure 1. Figure 2 shows example annotation pages within the annotation platform and the set of questions asked to the evaluators in case they find the provided LLM generated answer to be hallucinated (Figure 2(a)) or correct (Figure 2(b)).